### v2

# IEPile: Unearthing Large-Scale Schema-Based Information Extraction Corpus

Abstract
Large Language Models (LLMs) demonstrate remarkable potential across various domains; however, they exhibit a significant performance gap in Information Extraction (IE). Note that high-quality instruction data is the vital key for enhancing the specific capabilities of LLMs, while current IE datasets tend to be small in scale, fragmented, and lack standardized schema. To this end, we introduce IEPile, a comprehensive bilingual (English and Chinese) IE instruction corpus, which contains approximately 0.32B tokens. We construct IEPile by collecting and cleaning 33 existing IE datasets, and introduce schema-based instruction generation to unearth a large-scale corpus. Experimental results on LLaMA, Baichuan and Qwen demonstrate that using IEPile can enhance the performance of LLMs for IE, especially the zero-shot generalization. We open-source the resource and pre-trained models, hoping to provide valuable support to the NLP community.
IEPile: Unearthing Large-Scale Schema-Based
Information Extraction Corpus
Honghao ${}^{\spadesuit}$${}^{\diamondsuit}$thanks:     Equal Contribution., Lin ${}^{\clubsuit}$${}^{\diamondsuit}$11footnotemark: 1, Hongbin ${}^{\spadesuit}$, Ningyu ${}^{\spadesuit}$${}^{\diamondsuit}$thanks:     Corresponding Author. Mengshu ${}^{\clubsuit}$${}^{\diamondsuit}$, Lei ${}^{\clubsuit}$${}^{\diamondsuit}$, Huajun ${}^{\spadesuit}$${}^{\diamondsuit}$22footnotemark: 2 ${}^{\spadesuit}$ Zhejiang University ${}^{\diamondsuit}$ Zhejiang University - Ant Group Joint Laboratory of Knowledge Graph ${}^{\clubsuit}$ Ant Group {guihonghao,zhangningyu,huajunsir}@zju.edu.cn https://github.com/zjunlp/IEPile
1 Introduction
Large Language Models (LLMs) have achieved significant breakthroughs in multiple Natural Language Processing (NLP) tasks Du et al. (2022); Touvron et al. (2023b); Jiang et al. (2023); Zhao et al. (2023); Pu et al. (2023); Yang et al. (2024); Wu et al. (2023); Wang et al. (2023c); Fei et al. (2024). However, recent studies  (Li et al., 2023a; Ma et al., 2023; Xu et al., 2023; Wadhwa et al., 2023; Wan et al., 2023; Gao et al., 2023; Li et al., 2023b; Jiao et al., 2023; Huang et al., 2023; Wang et al., 2024) indicate a significant performance gap in the task of Information Extraction (IE) when utilizing LLMs.  (Lee et al., 2022a; Gao et al., 2023) further illustrate that the major reason may lie in limited high-quality, large-scale data corpus. Concretely, most IE datasets are often limited in size, scattered in distribution, and lack standardization in schema [^1] We refer to the schema as pre-defined types of entities, relations, events (arguments and roles), etc..
Faced with these limitations, there is an urgent need to collect instruction data in a unified and automated manner to build a high-quality, large-scale IE corpus. To this end, we collect and clean various existing IE datasets to obtain a comprehensive bilingual IE instruction dataset named IEPile [^2] IEPile adhere to the CC BY-NC-SA 4.0 license except for ACE2005 which adheres to the LDC User Agreement.. During the corpus construction, we find existing methods for constructing IE instruction data suffer from two issues for generalizable IE: 1) Schema Query Disparity: There may be inconsistency in the number of schema queries within instruction between training and evaluation which can harm model generalization; 2) Semantic Confusion: The co-occurrence of semantically similar schemas within instructions may confuse the model. Thus, we introduce a schema-based instruction generation strategy. We first construct a hard negative schema dictionary to promote the more frequent occurrence of semantically similar schema in instructions. Then, we introduce batched instruction generation, dynamically limiting the number of schemas queried in each instruction to $split\_num$, which not only addresses the issue of performance degradation due to inconsistent numbers of schema queries during training and evaluation, but also enhances the robustness when dealing with semantically confusing schema. Finally, we obtain IEPile which contains approximately 0.32B tokens.
By fine-tuning the Baichuan (Yang et al., 2023), LLaMA (Touvron et al., 2023b), and Qwen (Bai et al., 2023), we show that LLMs with IEPile can yield better zero-shot performance than baselines. This achievement not only verifies the effectiveness of the IEPile dataset but also provides a framework for creating IE datasets in other domains.
Figure 1: An overview of the construction of IEPile, including Data Collection and Cleaning, as well as Schema-Based Instruction Generation (Hard Negative Schema Construction and Batched Instruction Generation).
2 IEPile
In this section, we introduce the construction of IEPile and provide details in Appendix  B.
2.1 Data Collection and Cleaning
To broadly cover various domains and meet the practical demands, we collect datasets necessary for IE from multiple data sources. Our corpus mainly involves bilingual data (Chinese and English) and focuses on three main categories of IE tasks: Named Entity Recognition (NER), Relation Extraction (RE), and Event Extraction (EE). In total, we gather 26 English datasets and 7 Chinese datasets. We also employ standardization procedures to maintain data quality and format uniformity, involving format unification, instance deduplication, and the exclusion of low-quality data.
2.2 Schema-Based Instruction Generation
We concentrate on instruction-based information extraction (IE), which relies on three elements to compose an instruction: 1) Task Description, a template utilized to distinguish between different IE tasks; 2) Input Text, the source text that the model is expected to extract information; and 3) Schema sequence, which defines the information that the model is supposed to extract, including entity types, relations, events, etc. Among these, the schema sequence is critical as it reflects the specific extraction requirements and is dynamically variable. Therefore, constructing the schema sequence in an instruction is of utmost importance.
Positive and Negative Schema Mechanism in Instructions.
Firstly, we define schemas that actually exist within the input text as positive schemas and those that do not appear as negative schemas. As illustrated in Figure 1, the “location contains” present in the annotation is a positive schema, while all other schemas from the predefined label set $L$ are negative schemas. Traditional IE frameworks, which are treated as sequence labeling tasks, take text as input and produce a label for each token as output, without involving the concept of positive or negative schemas within the model’s input. However, in the era of generative IE, represented by models like UIE  (Lu et al., 2022a), introduce the concept of integrating a schema sequence (refers to as Structural Schema Instructor, or SSI) in the model’s input to guide its output, restricting the range of output to the SSI. The concept necessitates including the entire predefined label set of a dataset as the SSI to guide the model’s output during inference. Consequently, if the SSI during the training contains only positive schemas, the model will tend to generate corresponding answers for every label within the SSI during inference. Therefore, to make the model explicitly reject generating outputs for negative schemas, it is necessary to incorporate negative schemas into the SSI.
In this paper, the schema sequence included in the instructions follows the concept of SSI. However, we observe that existing research  (Wang et al., 2023b; Xiao et al., 2023) tends to adopt a rather crude schema processing strategy when constructing instructions, meaning that all schemas within a predefined label set are used to build the instructions. This approach potentially entails two significant issues: 1) Inconsistency in the number of schema queries within instruction between training and evaluation. For example, the model’s performance will decrease if it is trained on about 20 schema queries but tested with either 10 or 30, even if the training and evaluation schemas are similar in content. 2) Inadequate differentiation among schemas in the instructions. For example, semantically similar schemas like “layoffs”, “depart” and “dismissals”, may present co-occurrence ambiguities that could confuse the LLMs. Such schemas should co-occur more frequently within the instruction. Therefore, we introduce: 1) Hard Negative Schema Construction; and 2) Batched Instruction Generation. Detailed information can be found in Figure 1 and Algorithm 1.
Hard Negative Schema Construction.
As illustrated in Figure 1, assume that dataset $\mathcal{D}$ possesses a predefined label set $L$. For a given text $S$, the schemas present in its annotation constitute the positive schema set $Pos\_L$, while others form the negative schema set $Neg\_L$. In our analysis, we discover that the primary cause of model mistakes stems from the semantic ambiguity of the schema. In traditional approaches, the $Neg\_L$ is simply defined as $L-Pos\_L$. However, they overlook a critical aspect: it is important to pay special attention to negative schemas that are semantically close to positive schemas. Inspired by the theory of contrastive learning, we construct a hard negative schema dictionary $\mathcal{K}$, where each key represents a unique schema and the associated value is a collection of schemas that are semantically similar to the key schema. Based on this, we define the hard negative schema set as $Hard\_L=\mathcal{K}[Pos\_L]$, and the other negative schema set as $Other\_L=L-Pos\_L-Hard\_L$. The final $Neg\_L$ is constituted by $Hard\_L$ and a small subset of $Other\_L$. Through this strategy, we not only present semantically similar schemas more frequently within the instruction but also reduce the number of training instances without sacrificing model performance.
Batched Instruction Generation.
Subsequently, we obtain the final schema set $L^{\prime}=Pos\_L+Neg\_L$. We employ a batched instruction generation method, dynamically limiting the number of schemas inquired in each instruction to the number of $split\_num$, which ranges between 4 and 6. Therefore, $L^{\prime}$ will be divided into $|L^{\prime}|/split\_num$ batches for querying, with each batch querying $split\_num$ schemas. Consequently, even if the number of schemas inquired during the evaluation phase differs from that of training, the batched mechanism allows us to distribute the inquiries across $split\_num$ schemas, thereby mitigating the decline in generalization performance.
Figure 2: Distribution of different tasks, domains, and datasets within IEPile.
2.3 Data Statistics
Based on the aforementioned methods, we obtain IEPile, which includes roughly 2 million instruction entries and approximately 0.32B tokens (utilizing the Baichuan2 tokenizer). Figure 2 displays the distribution of domains and datasets within IEPile, including 33 datasets spanning multiple domains such as general, news, finance, and biomedical. Additionally, Table B.2 provides examples of instructions and outputs for 3 different tasks within the IEPile dataset.
Method NER RE EE
CrossNER FewRel Wiki-ZSL Avg WikiEvents RAMS CrudeOil News Avg
LLaMA2 34.82 6.53 9.43 7.98 0.00 0.00 0.00 0.00
Baichuan2 38.93 5.94 4.15 5.05 0.00 0.00 0.00 0.00
Qwen1.5 50.13 7.82 6.94 7.38 0.00 0.00 0.00 0.00
Mistral 42.83 6.84 5.10 5.97 0.00 0.00 0.00 0.00
ChatGPT 58.37 9.96 13.14 11.55 2.95 8.35 1.41 4.24
GPT-4 58.49 22.43 23.76 23.10 5.24 10.14 26.13 13.84
UIE 38.37 - - - 5.12 9.25 6.45 6.94
InstructUIE 49.36 39.55 35.20 37.38 11.64 24.27 23.26 19.72
YAYI-UIE 50.39 36.09 41.07 38.58 10.97 18.87 12.45 14.10
Baichuan2-IEPile 55.55 41.28 37.61 39.45 9.12 20.19 36.61 21.97
LLaMA2-IEPile 56.50 37.14 36.18 36.66 13.93 23.62 33.87 23.81
Qwen1.5-IEPile 57.90 40.92 38.49 39.71 11.38 21.26 30.69 21.11
Table 1: Zero-shot performance on English datasets. UIE necessitates predefined entity types; given that such information is not provided by the FewRel and Wiki-ZSL datasets, we are unable to evaluate UIE’s performance on these datasets. For the task of event extraction, we only present the results of event detection in the main text.
Method NER RE EE
Boson Weibo Avg SKE2020 COAE2016 IPRE Avg FewFC CCF Law Avg
LLaMA2 8.19 2.43 5.31 0.50 3.11 0.31 1.31 0.23 0.08 0.16
Baichuan2 27.39 7.62 17.51 7.23 11.65 1.45 6.78 11.82 2.73 7.28
Qwen1.5 26.49 25.34 25.92 7.69 11.97 2.16 7.27 11.47 3.25 7.36
Mistral 29.13 10.02 19.58 6.84 5.24 0.82 4.30 4.69 0.23 2.46
ChatGPT 38.53 29.30 33.92 24.47 19.31 6.73 16.84 16.15 0.00 8.08
GPT-4 48.15 29.80 38.98 56.77 41.15 18.15 38.69 74.25 42.12 58.19
YAYI-UIE 49.25 36.46 42.86 70.80 19.97 22.97 37.91 81.28 12.87 47.08
Baichuan2-IEPile 55.77 38.03 46.90 72.50 47.43 29.76 49.90 83.59 63.53 73.56
LLaMA2-IEPile 54.45 34.97 44.71 72.18 46.70 28.55 49.14 70.10 59.90 65.00
Qwen1.5-IEPile 63.08 37.50 50.29 72.29 50.70 30.55 51.18 78.77 61.43 70.10
Table 2: Zero-shot performance on Chinese datasets. Since UIE and InstructUIE do not train with Chinese data, we do not report performance of these two models on Chinese datasets.
3 Experiments
Based on IEPile, we fine-tune (Lora) models such as Baichuan2, LLaMA2, and Qwen1.5, subsequently comparing their zero-shot generalization capabilities against a range of baseline models. Results of the full supervision evaluation, and training details are in Appendix C.
3.1 Experimental Settings
Evaluation Metrics: We employ span-based Micro-F1 as the metric for measuring model performance. Baselines: We select a range of strong models for comparative analysis, which include UIE  (Lu et al., 2022a), LLaMA2-13B-Chat, Baichuan2-13B-Chat, Mistral-7B-Instruct-v0.2  (Jiang et al., 2023), Qwen1.5-14B-Chat, ChatGPT  (Ouyang et al., 2022), GPT-4  (OpenAI, 2023), InstructUIE  (Wang et al., 2023b), and YAYI-UIE  (Xiao et al., 2023). Zero-shot Benchmark: We collect 13 datasets that are not present in the training set.
3.2 Main Results
In Tables 1 and  2, we report the zero-shot performance across three tasks: NER, RE, and EE for two languages. Overall, after training with the IEPile, the models achieve better results in the majority of tasks. We believe the success is due to the hard negative schema construction and batched instruction generation strategy, which can mitigate the train-eval mismatch and semantic ambiguity for the diverse schema. We also observe that Baichuan2-IEPile, LLaMA2-IEPile and Qwen1.5-IEPile are slightly behind GPT-4 in English NER. We hypothesize that the marginal gap may be attributed to GPT-4’s exposure to a vast corpus of similar data during its training. Moreover, it is essential to note that InstructUIE focuses on English data while IEPile incorporates both English and Chinese data. This disparity in data may influence the capability of the model in English, potentially reducing the performance.
Inconsistency in the Number of Schema Queries Hurt Generalization.
We investigate the impact on model performance when different numbers of schema queries are used during the training and evaluation. We train the Baichuan2 using full-schema instructions on 3 datasets: Ontonotes (18 schemas), DuIE2.0 (49 schemas), and ACE2005 (33 schemas). For the evaluation, we test the model using two strategies: one with the full set of schema queries and another with a fixed set of 10 schema queries. The results depicted in Figure 3(a) indicate that the mismatch in the number of schema queries during evaluation significantly reduces the model’s performance. Further analysis of the model’s outputs reveals that the model always tends to generate outputs for each inquiry. We hypothesize that the number of schema queries is one of the key factors affecting the generalization ability. The model needs to first adapt to the number of schema inquiries that are rare during the training and then adapt to the unseen schema.
Figure 3: (a) When there is an inconsistency in the number of schema inquiries during the training and evaluation, the performance of the model significantly decreases. (b) The impact of removing the hard negative schema dictionary on the performance of the model.
3.3 Analysis
Inadequate Differentiation Among Schemas Lead to Semantic Similar Confusion.
We also evaluate the impact of removing the “Hard Negative Schema Dictionary” on the performance of Baichuan2-IEPile, with particular attention to schemas that are hard to differentiate. According to the results in Figure 3(b), we notice that the hard negative schema dictionary plays a relatively limited role in the NER task, which may be due to the clear boundaries inherent to entity recognition. However, the utilization of the hard negative schema dictionary notably enhances model performance in the DuIE2.0 and DuEE1.0 datasets. We observe that semantically similar and easily confused schemas frequently appeared in the model’s outputs, such as predicting “dismissal” and “resignation” in the event of “layoff”. Therefore, processing instructions that are semantically prone to confusion poses significant challenges, and the hard negative schema dictionary plays a crucial role in bolstering model robustness and improving the accuracy of predictions.
4 Conclusion and Future Work
In this paper, we introduce IEPile, by collecting and cleaning existing IE datasets and utilizing a schema-based instruction generation strategy. In the future, we will continue to maintain the corpus and try to integrate new resources including open-domain IE, and document-level IE.
Limitations
From the data perspective, our study primarily focuses on schema-based IE, which limits our ability to generalize to human instructions that do not follow our specific format requirements. Additionally, we do not explore the field of Open Information Extraction (Open IE); however, if we remove schema constraints, our dataset would be suitable for Open IE scenarios. Besides, IEPile is confined to data in English and Chinese, and in the future, we hope to include data in more languages. From the model’s perspective, our research evaluates only two models, Baichuan and LLaMA, along with a few baselines due to the computation resources. IEPile can be applied to any other LLMs such as ChatGLM Du et al. (2022) and Gemma [^3] https://huggingface.co/google/gemma-7b..
Ethical Considerations
In this paper, we strictly adhered to the standards and principles of ethics. All data collected are sourced from publicly available materials, ensuring the transparency and legality of the research. We conduct a thorough review of the data used, verifying the legitimacy of their sources and compliance with their usage, thus avoiding any infringement on personal privacy or involvement with unauthorized information.
Appendix A Related Work
A.1 Information Extraction Datasets
Large-scale pre-trained corpora are crucial for the effectiveness of LLMs, providing a wealth of knowledge and a foundation for language comprehension, while the annotated data for information extraction (IE) also holds its importance. Although the field of IE has accumulated a considerable amount of annotated data (Walker et al., 2006; Riedel et al., 2010; Sang and Meulder, 2003; Luan et al., 2018), these datasets are often limited in size, scattered in distribution, and lack standardization in schema. Faced with these limitations, there is an urgent need for generating instruction data through unified and automated methods to fill the gap of the current lack of centralized, large-scale IE instruction tuning datasets. We focus on instruction-based IE scenarios and, by collecting and cleaning existing IE data, build a large, schema-rich IE instruction dataset called IEPile. IEPile is designed to enhance the adaptability and processing capabilities of LLMs for different IE tasks, simultaneously strengthening their generalization skills to extract from new domains and schemas.
A.2 Information Extraction Models
Recently, LLMs (Brown et al., 2020; Ouyang et al., 2022; Touvron et al., 2023a, b) demonstrate their exceptional versatility and generalization capabilities across a variety of downstream tasks  (Vilar et al., 2023; Hegselmann et al., 2023). Particularly in the domain of IE, these models have the potential to tackle many challenges previously encountered in research  (Zheng et al., 2017; Li et al., 2020a; Paolini et al., 2021; Lu et al., 2022b; Lou et al., 2023; Chen et al., 2022b, 2024), such as adaptability issues when dealing with unseen labels during training. Some studies  (Wei et al., 2023; Wang et al., 2023a; Xie et al., 2023) make significant performance gains in low-resource settings by designing prompt-based frameworks and leveraging models like ChatGPT for in-context learning. Moreover, research efforts such as InstructUIE  (Wang et al., 2023b), PIVOINE  (Lu et al., 2023), and YAYI-UIE  (Xiao et al., 2023), which employ instruction-tuning of open-source LLMs, also achieve notable successes on IE. Additional research explore areas such as prompt learning  (Zhang et al., 2023a), guidelines  (Sainz et al., 2023) and synthetic dataset  (Amalvy et al., 2023). Despite these advancements, current models fine-tuned with an instruction approach face a major challenge: the coarse schema handling strategies in constructing instructions can lead to a mismatch in the number of schema queries between the training and evaluation phases, thus weakening the model’s generalizability.
Appendix B Construction Details of IEPile
B.1 Data Collection and Clean
Data Collection
To comprehensively cover various domains and meet the practical demands of IE, we collect IE datasets from multiple sources. IEPile dataset mainly involves bilingual data (Chinese and English) and focuses on three main categories of IE tasks: Named Entity Recognition (NER), Relation Extraction (RE), and Event Extraction (EE). The English data mainly comes from the benchmark dataset IEINSTRUCTIONS  (Wang et al., 2023b), while the Chinese data is similar to the Chinese datasets mentioned in the YAYI-UIE  (Xiao et al., 2023). It should be noted that our Chinese dataset collection is conducted concurrently with the aforementioned research.
Specifically, the NER datasets include fifteen English datasets such as ACE2005  (Walker et al., 2006), AnatEM  (Pyysalo and Ananiadou, 2014), BC2GM  (Kocaman and Talby, 2020), BC4CHEMD  (Kocaman and Talby, 2020), BC5CDR  (Zhang et al., 2023b), CoNLL2003  (Sang and Meulder, 2003), FabNER  (Kumar and Starly, 2022), FindVehicle  (Guan et al., 2023), GENIA-Ent  (Kim et al., 2003), HarveyNER  (Chen et al., 2022a), MIT Movie  (Liu et al., 2013), MIT Restaurant  (Liu et al., 2013), MultiNERD  (Tedeschi and Navigli, 2022), NCBI-Disease  (Dogan et al., 2014), Ontonotes  (Pradhan and Xue, 2009), and three Chinese datasets including MSRA  (Levow, 2006), Resume NER  (Zhang and Yang, 2018), CLUE NER  (Xu et al., 2020). The RE task encompasses eight English datasets including ADE Corpus  (Gurulingappa et al., 2012), CoNLL2004  (Carreras and Màrquez, 2004), GIDS  (Jat et al., 2017), KBP37  (Zhang and Wang, 2015), NYT  (Riedel et al., 2010), NYT11-HRL  (Takanobu et al., 2019), SciERC  (Luan et al., 2018), Semeval-RE  (Hendrickx et al., 2010), and two Chinese datasets, CMeIE  (Luan et al., 2018), DuIE2.0  (Hendrickx et al., 2010). The EE task covers three English datasets: ACE2005  (Walker et al., 2006), CASIE  (Satyapanich et al., 2020), PHEE  (Sun et al., 2022), and two Chinese datasets, DuEE1.0  (Satyapanich et al., 2020), DuEE-fin  (Sun et al., 2022). These datasets span various domains such as general, medical, financial, and more. For more detailed statistical information, please refer to Tables 9,  10 and  11.
Data Cleaning
During the data cleaning process, we address each dataset individually. Firstly, we calculate the text overlap within each dataset’s training, validation, and test sets. If a text is discovered to have multiple occurrences within the same file accompanied by inconsistent annotations, we exclude all corresponding instances from the dataset. Secondly, we compare the text overlap between training, validation, and test sets. If texts from the test set appear previously in the training or validation sets, we would exclude these instances from the training and validation sets. Moreover, we formulate three heuristic rules to eliminate low-quality and meaningless data:
1) Non-alphabetic characters comprising more than 80% of the text;
2) Text length under five characters without any labels;
3) A high prevalence of stopwords such as ‘the,’ ‘to,’ ‘of,’ etc., exceeding 80%.
We believe that the aforementioned cleaning measures will positively affect model training and enhance its performance. Moreover, our efforts unify data formats across various tasks and conduct a thorough audit of each dataset, creating detailed data records that include the volume of data, domains, schemas, and other information. Figure 4 is an example of a data record for Ontonotes.
Figure 4: An exemplar of data records for OntoNotes: the domain, the number and details of schemas, the total volume of data, the $split\_num$, the number of instructions produced using our method, along with the distribution of split count within the interval [($split\_num$ / 2), ($split\_num$ + $split\_num$ / 2)].
B.2 Schema-Based Instruction Generation
Hard Negative Schema Construction.
As illustrated in Figure 1, assume that dataset $\mathcal{D}$ possesses a predefined label set $L$. For a given text $S$, the schemas present in its annotation constitute the positive schema set $Pos\_L$, while others form the negative schema set $Neg\_L$. Inspired by the theory of contrastive learning, we construct a hard negative schema dictionary $\mathcal{K}$, where each key represents a unique schema and the associated value is a collection of schemas that are semantically similar to the key schema. Consequently, the set of hard negative schema, $Hard\_L$, is defined as $\mathcal{K}[Pos\_L]$. However, if $Neg\_L$ is composed solely of $Hard\_L$, it would lack a sufficient number of negative instances for the model to learn effectively. Therefore, we define another set of negative schemas, $Other\_L=L-Hard\_L-Pos\_L$. Ultimately, the $Neg\_L$ is composed of $Hard\_L$ and a small number of $Other\_L$ (roughly $split\_num$). The rationale behind the development of these hard negatives is two-fold: firstly, to induce a more frequent co-occur of semantically similar schemas within the instructions, and secondly, to reduce the volume of training instances without sacrificing the model’s performance. Consider a scenario involving 48 schemas paired with a $split\_num$ of 4. Conventionally, this setup would necessitate the generation of 12 distinct instructions. Yet, with the inclusion of hard negatives, this number can be reduced to as few as 3.
Algorithm 1 Schema-Based Instruction Generation
0:  Text $S$, Full label set $L$, Positive label set $Pos\_L$, Number of schemas to split $split\_num$
0:  Negative label set $Neg\_L$, Set of instructions $Instructions$ Step 1: Initialize Hard Negative Schema Dictionary $\mathcal{K}$
  for all schema in $L$ do
     $\mathcal{K}[schema]\leftarrow\text{SEMANTIC-CLOSE}(schema,L)$
  end for
  Step 2: Obtain Hard Negative Schemas
  $Hard\_L\leftarrow\emptyset$
  for all schema in $Pos\_L$ do
     $Hard\_L\leftarrow Hard\_L\cup\mathcal{K}[schema]$
  end for
  $Other\_L\leftarrow L-Pos\_L-Hard\_L$
  $Other\_L\leftarrow\text{RANDOM-SELECT}(Other\_L,split\_num)$
  $Neg\_L\leftarrow Hard\_L\cup Other\_L$
  $L^{\prime}=Neg\_L\cup Pos\_L$
  Shuffle $L^{\prime}$ to get a randomized sequence
  Step 3: Batched Instruction Generation
  $Instructions\leftarrow[]$
  $num\_batches\leftarrow\lceil\frac{|L^{\prime}|}{split\_num}\rceil$
  for $i\leftarrow 1$ to $num\_batches$ do
     $Batch\leftarrow\text{SEQUENTIAL-SELECT}(L^{\prime},split\_num,i)$
     $Instructions\leftarrow Instructions\cup\text{GENERATE-INSTRUCTION}(Batch)$
  end for
Batched Instruction Generation.
Subsequently, we obtain the final schema set $L^{\prime}=Pos\_L+Neg\_L$. During the instruction generation phase, the role of schemas is critically vital, as it reflects the specific extraction requirements and is dynamically variable. Traditional practices typically integrate the full schema set into the instruction. However, in this study, we employ a batched instruction generation method, dynamically limiting the number of schemas inquired in each instruction to the number of $split\_num$, which ranges between 4 to 6. Therefore, $L^{\prime}$ will be divided into $|L^{\prime}|/split\_num$ batches for querying, with each batch querying $split\_num$ schemas. Consequently, even if the number of schemas inquired during the evaluation phase differs from that of training, the batched mechanism allows us to distribute the inquiries across $split\_num$ schemas, thereby mitigating the decline in generalization performance.
Selection of split_num.
In determining the range for $split\_num$, we combine empirical experimental results with an in-depth analysis of dataset characteristics. For a dataset containing N different labels, the theoretical value of $split\_num$ should fall within the interval [1, N]. Given that we deal with datasets of varying numbers of labels, we aim to identify a $split\_num$ value that could be shared by as many datasets as possible, thereby ensuring that this number is a common divisor for the label numbers of most datasets. For instance, for Named Entity Recognition datasets, we set $split\_num$ to 6; for Relation Extraction and Event Extraction datasets, we establish $split_{n}um$ at 4. We also observe that when $split\_num$ is 1, the ratio of positive to negative samples significantly impacts model performance, and the corresponding number of training samples becomes vast, affecting efficiency adversely. More crucially, we believe that enumerating multiple schemas in instructions aids the model in more effectively learning to distinguish and identify various schemas, thereby enhancing model performance.
Moreover, to enhance model robustness and its clear understanding of the dynamically changing schema sequences in instructions, we set the actual number of schema splits within a dynamic range of [$split\_num$ // 2, $split\_num$ + $split\_num$ // 2]. Specifically, if the number of schemas in the last batch is less than half of $split\_num$, it is merged with the previous batch; it is only treated as a separate batch when its number exceeds half of $split\_num$.
Task Instruction & Output
NER
1{
2    "instruction": "You are an expert in named entity recognition. Please extract entities that match the schema definition from the input. Return an empty list if the entity type does not exist. Please respond in the format of a JSON string.",
3    "schema": ["location", "else", "organization", "person"],
4    "input": "The objective of the Basic Course on War is to provide for combatants of the EPR basic military knowledge for the armed conflict against the police and military apparatus of the bourgeoisie."
5}
6output = {
7    "location": [],
8    "else": [],
9    "organization": ["EPR"],
10    "person": []
11}
RE
1{
2    "instruction": "You are an expert in relationship extraction. Please extract relationship triples that match the schema definition from the input. Return an empty list for relationships that do not exist. Please respond in the format of a JSON string.",
3    "schema": ["place of birth", "country capital", "country of administrative divisions", "company"],
4    "input": "Born on May 1 , 1927 , in Brichevo , Bessarabia in the present-day Republic of Moldova , Mr. Bertini emigrated to Palestine with his family as a child and pursued musical studies there , in Milan , and in Paris , where he worked with Nadia Boulanger and Arthur Honegger."
5}
6output = {
7    "place of birth": [{"head": "Mr. Bertini", "tail": "Paris"}],
8    "country capital": [],
9    "country of administrative divisions": [],
10    "company": []
11}
EE
1{
2    "instruction": "You are an expert in event extraction. Please extract events from the input that conform to the schema definition. Return an empty list for events that do not exist, and return NAN for arguments that do not exist. If an argument has multiple values, please return a list. Respond in the format of a JSON string.",
3    "schema": [{"event_type": "pardon", "trigger": true, "arguments": ["defendant"]}, {"event_type": "extradite", "trigger": true, "arguments": ["person", "agent", "destination", "origin"]}, {"event_type": "sue", "trigger": true, "arguments": ["place", "plaintiff"]}, {"event_type": "start position", "trigger": true, "arguments": ["person", "entity", "place"]}],
4    "input": "Ethical and legal issues in hiring Marinello"
5}
6output = {
7    "pardon": [],
8    "extradite": [],
9    "sue": [],
10    "start position": [{"trigger": "hiring", "arguments": {"person": "Marinello", "entity": "NAN", "place": "NAN"}}]
11}
Table 3: Instructions and outputs for 3 tasks: Named Entity Recognition (NER), Relation Extraction (RE), and Event Extraction (EE). The instruction and output formats for IEPile adopt a structure similar to JSON strings.
Instruction Format
The instruction format of IEPile adopts a structure akin to JSON strings, essentially constituting a dictionary-type string. This structure is comprised of three main components: (1) “instruction”, which is the task description outlining the objective of the instruction’s execution; (2) “schema”, a list of labels that need to be extracted, specifying the key fields of the information to be extracted; (3) “input”, the source text from which information is to be extracted. Examples of instructions corresponding to various tasks can be found in Table B.2.
B.3 Data Statistics
Based on the aforementioned methodologies, we construct a high-quality information extraction instruction dataset known as IEPile. This dataset contains approximately two million instances, each comprising two fields: “instruction” and “output”, which are formatted for direct use in the supervised fine-tuning of models. In terms of storage, IEPile occupies around 3GB of disk space, encompassing roughly 0.32B tokens (utilizing the BaiChuan2 tokenizer).
Appendix C Experiments
C.1 Experimental Settings
Evaluation Metrics
We employ span-based Micro-F1 as the primary metric for measuring model performance. For the NER task, the model is required to accurately identify the boundaries of entities and their corresponding types. For the RE task, the model must precisely determine the subject and object entities within a relation, as well as the type of relation between them. As for the EE task, we match the event triggers, denoted as Trigger, and the arguments, referred to as Argument, independently.
Baselines
For evaluating zero-shot generalization capabilities, we select a range of strong models for comparative analysis:
  • UIE  (Lu et al., 2022a): is a unified text-to-structure generation framework that can model various information extraction (IE) tasks generically.
  • LLaMA2  (Touvron et al., 2023b): is a series of LLMs ranging from 7 billion to 70 billion parameters.
  • Baichuan2  (Yang et al., 2023): is a collection of multilingual LLMs containing 7 billion and 13 billion parameters.
  • Qwen1.5  (Bai et al., 2023): is a comprehensive language model series that encompasses distinct models with varying parameter counts.
  • Mistral  (Jiang et al., 2023): is a 7-billion-parameter LLM.
  • ChatGPT  (Ouyang et al., 2022): also known as GPT-3.5-turbo, represents the most advanced artificial intelligence language model with chat optimization capabilities to date.
  • InstructUIE  (Wang et al., 2023b): a unified IE framework based on multi-task instruction tuning.
  • YAYI-UIE  (Xiao et al., 2023): is an end-to-end, chat-enhanced, universal information extraction framework that supports both Chinese and English, fine-tuned with instructional prompts for generalized information.
Zero-shot Dataset
To ensure the validity of the zero-shot evaluation and prevent result bias due to data similarity, we select datasets primarily derived from news and biomedical fields as our training sets. This selection is intended to train the model’s capability for instruction following and schema-based extraction. For the evaluation data, we adopt the 13 cross-domain datasets recommended in IEINSTRUCTIONS and YAYI-UIE, which include: for Named Entity Recognition (NER) tasks, we use the CrossNER  (Liu et al., 2021), Weibo NER  (Peng and Dredze, 2015), and Boson [^4] https://github.com/InsaneLife/; in Relation Extraction (RE) tasks, we involve FewRel  (Han et al., 2018), Wiki-ZSL  (Chen and Li, 2021), COAE2016555https://github.com/Sewens/COAE2016, IPRE  (Wang et al., 2019), and SKE2020666https://aistudio.baidu.com/datasetdetail/177191; and for Event Extraction (EE), we include RAMS  (Ebner et al., 2020), WikiEvents  (Li et al., 2021), CrudeOilNews  (Lee et al., 2022b), FewFC  (Zhou et al., 2021), and CCF law  [^7] https://aistudio.baidu.com/projectdetail/4201483. These datasets cover a wide range of fields including literature, music, law, and oil news. It is noteworthy that these evaluation data sets are not used during the training, ensuring that our evaluation accurately reflects the model’s generalization and adaptation capabilities for unseen domains and unseen schema data in zero-shot information extraction.
Hyperparameter Value
Number of Epochs 5
Learning Rate 5e-5
Batch Size 20
Accumulate 4
Optimizer Adam
Lora_r 64
Lora_alpha 64
Lora_dropout 0.05
Table 4: Training Hyperparameters
Method EN CH
WikiEvents RAMS
CrudeOil
News
Avg FewFC
CCF
Law
Avg
Trigger LLaMA2 0.00 0.00 0.00 0.00 0.23 0.08 0.16
Baichuan2 0.00 0.00 0.00 0.00 11.82 2.73 7.28
Qwen1.5 0.00 0.00 0.00 0.00 11.47 3.25 7.36
Mistral 0.00 0.00 0.00 0.00 4.69 0.23 2.46
ChatGPT 2.95 8.35 1.41 4.24 16.15 0.00 8.08
GPT4.0 5.24 10.14 26.13 13.84 74.25 42.12 58.19
UIE 5.12 9.25 6.45 6.94 - - -
InstructUIE 11.64 24.27 23.26 19.72 - - -
YAYI-UIE 10.97 18.87 12.45 14.10 81.28 12.87 47.08
Baichuan2-IEPile 9.12 20.19 36.61 21.97 83.59 63.53 73.56
LLaMA2-IEPile 13.93 23.62 33.87 23.81 70.10 59.90 65.00
Qwen1.5-IEPile 11.38 21.26 30.69 21.11 78.77 61.43 70.10
Argument LLaMA2 0.00 0.00 0.00 0.00 0.00 0.06 0.03
Baichuan2 0.79 1.81 0.48 1.03 6.91 13.04 9.98
Qwen1.5 0.64 2.31 0.74 1.23 6.37 14.48 10.43
Mistral 0.24 0.65 0.16 0.35 7.43 6.60 7.02
ChatGPT 2.07 2.21 8.60 4.29 44.40 44.57 44.49
GPT4.0 3.35 7.35 17.25 9.32 48.05 47.49 47.77
UIE 1.78 2.14 8.95 4.29 - - -
InstructUIE 5.88 6.21 21.78 11.29 - - -
YAYI-UIE 5.11 8.21 19.74 11.02 63.06 59.42 61.24
Baichuan2-IEPile 7.64 10.42 20.40 12.82 57.93 65.43 61.68
LLaMA2-IEPile 12.55 11.30 18.47 14.11 43.26 35.71 39.49
Qwen1.5-IEPile 11.93 10.57 20.22 14.24 59.49 58.86 59.18
Table 5: Zero-shot performance on Event Extraction (EE) task. Within each column, shadow and shadow represent the top 2 results.
C.2 Zero-shot performance on Event Extraction
As illustrated in Table  5, the model trained with IEPile exhibits outstanding performance in zero-shot event extraction (EE) tasks, surpassing other competing models. Notably, in the Chinese EE task, the LLaMA2-IEPile model’s performance is slightly inferior to YAYI-UIE’s, revealing LLaMA2’s limitations in processing Chinese data. However, in the English EE task, LLaMA2-IEPile’s performance is significantly superior to that of similar models. This contrast highlights the potential influence of language type on model performance.
C.3 Hyper-parameter
In our research, we select two pre-trained models, Baichuan2-13B-Chat and LLaMA2-13B-Chat, as the base models for our study. Specifically, we employ the LoRA  (Hu et al., 2022) technique and utilize 8 NVIDIA A800 GPUs to perform instruction tuning on our IEPile dataset. The training period for each model spans approximately three days. Detailed configurations of the hyperparameters during the fine-tuning process are presented in Table 4.
Dataset InstructUIE YAYI-UIE Baichuan2-IEPile LLaMA2-IEPile
ACE2005 86.66 81.78 81.86 81.14
AnatEM 90.89 76.54 87.21 86.90
BC2GM 85.16 82.05 80.73 83.07
BC4CHEMD 90.30 88.46 90.45 90.07
BC5CDR 89.59 83.67 88.07 88.01
CoNLL2003 92.94 96.77 92.49 92.98
FabNER 76.20 72.63 77.07 76.33
FindVehicle 89.47 98.47 98.49 97.91
GENIA-Ent 74.71 75.21 76.66 77.32
HarveyNER 88.79 69.57 67.70 62.64
MIT Movie 89.01 70.14 88.23 89.54
MIT Restaurant 82.55 79.38 79.85 81.30
MultiNERD 92.32 88.42 94.60 94.24
NCBI-Disease 90.23 87.29 85.26 87.59
Ontonotes 90.19 87.04 87.55 90.34
Avg 87.27 82.49 85.08 85.29
MSRA - 95.57 87.99 86.32
Resume NER - - 93.92 92.86
CLUE NER - - 80.19 76.57
Table 6: Overall supervision results on Named Entity Recognition (NER) datasets. Within each row, shadow and shadow represent the top 2 results.
Dataset InstructUIE YAYI-UIE Baichuan2-IEPile LLaMA2-IEPile
ADE Corpus 82.31 84.14 83.73 85.87
CoNLL2004 78.48 79.73 72.87 73.71
GIDS 81.98 72.36 74.71 74.13
KBP37 36.14 59.35 65.09 61.49
NYT 90.47 89.97 93.0 92.22
NYT11-HRL 56.06 57.53 53.19 54.86
SciERC 45.15 40.94 43.53 44.58
Semeval-RE 73.23 61.02 58.47 57.61
Avg 67.98 68.13 68.07 68.06
CMeIE - - 49.16 47.40
DuIE2.0 - 81.19 75.61 74.34
Table 7: Overall supervision results on Relation Extraction (RE) datasets. Within each row, shadow and shadow represent the top 2 results.
Dataset InstructUIE YAYI-UIE Baichuan2-IEPile LLaMA-IEPile
Trigger ACE2005 77.13 65.00 72.46 70.63
CASIE 67.80 63.00 60.07 61.27
PHEE 70.14 63.00 66.22 68.52
Avg 71.69 63.67 66.25 66.81
DuEE1.0 - 85.00 86.73 84.01
DuEE-fin - 82.50 83.54 79.00
Argument ACE2005 72.94 62.71 63.90 62.69
CASIE 63.53 64.23 56.07 56.78
PHEE 62.91 77.19 70.85 71.33
Avg 66.46 68.04 63.60 63.61
DuEE1.0 - 79.08 75.63 73.79
DuEE-fin - 70.02 79.34 73.08
Table 8: Overall supervision results on Event Extraction (EE) datasets. Within each row, shadow and shadow represent the top 2 results.
C.4 Supervision Results
Tables  6,  7, and  8 present our experimental results under a supervised learning setting on the training dataset. Specifically, it can be observed that after training on the IEPile, the model excels in Named Entity Recognition (NER), Relation Extraction (RE), and Event Detection (ED), ranking second across these tasks. The model’s performance is only slightly behind other reference models in the Event Argument Extraction. Additionally, we record the model’s performance in Chinese NER, RE, and EE tasks, where it demonstrates robust results. Notably, unlike fully parameter-fine-tuned models such as InstructUIE and YAYI-UIE, we employ the Low-Rank Adaptation (LoRA) technique for model optimization, which involves partial retraining. This difference in fine-tuning approach may have a certain impact on the results of supervision learning. In a comprehensive assessment, the IEPile-trained model showcases performance on par with other models in instruction-based information extraction (IE) tasks and significantly improves performance in zero-shot IE tasks compared to other models. This indicates the significant application prospects and potential of IEPile in the current field of IE.
C.5 Impact of Potential Dataset Bias on Model Performance and Generalization
During the IEPile project, we identify that potential biases introduced by the datasets used can affect the model’s performance and generalization capability. Firstly, biases in the definition of schemas within the datasets have a negative impact on model performance  (Huang et al., 2024). In the early stages of training, we observe instability in results due to mutual interference among multiple datasets that contain the same schemas but with differing definitions (such as wikiann, wikineural, polyglot-NER, and CoNLL2003). Consequently, in the later stages, we remove all such datasets except for CoNLL2003. Secondly, the model demonstrates good generalization when dealing with datasets having schemas similar to those in the training set. For instance, the model proficiently handles People Daily and ACE2004 datasets, which share similar schemas to the MASR and ACE2005 NER datasets in the training set, despite not being included in the training corpus. Lastly, the use of common, coarse-grained labels (such as “person” and “organization”) within the IEPile project lead the model, after training, to favor these coarse categories over fine-grained ones (such as “scientist” and “company”) when predicting instructions that included both levels of granularity.
Task Dataset Domain #Schemas #Train #Val #Test
NER-en AnatEM  (Pyysalo and Ananiadou, 2014) Biomedical 1 5667 2081 3758
BC2GM  (Kocaman and Talby, 2020) Biomedical 1 12392 2483 4977
BC4CHEMD (Kocaman and Talby, 2020) Biomedical 1 30488 30468 26204
NCBI-Disease  (Dogan et al., 2014) Biomedical 1 5432 923 940
BC5CDR (Zhang et al., 2023b) Biomedical 2 4545 4569 4788
HarveyNER  (Chen et al., 2022a) Social Media 4 3553 1270 1260
CoNLL2003  (Sang and Meulder, 2003) News 4 12613 3070 3184
GENIA  (Kim et al., 2003) Biomedical 5 14966 1657 1850
ACE2005  (Walker et al., 2006) News 7 7134 964 1050
MIT Restaurant  (Liu et al., 2013) Social Media 8 7658 - 1520
MIT Movie  (Liu et al., 2013) Social Media 12 9707 - 2441
FabNER  (Kumar and Starly, 2022) Scientific 12 9421 2179 2064
MultiNERD  (Tedeschi and Navigli, 2022) Wikipedia 16 130623 9994 9994
Ontonotes  (Pradhan and Xue, 2009) General 18 54994 7997 7782
FindVehicle  (Guan et al., 2023) Traffic 21 21547 - 20769
[…]
NER-zh MSRA NER  (Levow, 2006) News 3 40500 4500 3437
Resume NER  (Zhang and Yang, 2018) Resume 8 3799 463 476
CLUE NER  (Xu et al., 2020) News 10 9674 1074 1343
[…]
Table 9: Statistical data of Named Entity Recognition (NER) datasets, with an asterisk ($\dagger$) indicating a zero-shot evaluation set not included in the instruction tuning. CrossNER (Liu et al., 2021) is divided into five subsets for our statistical analysis. For the zero-shot evaluation sets, we only report their test set.
Task Dataset Domain #Schemas #Train #Val #Test
RE-en ADE Corpus  (Gurulingappa et al., 2012) Biomedical 1 3416 427 428
GIDS  (Jat et al., 2017) News 4 8525 1417 4307
CoNLL2004  (Carreras and Màrquez, 2004) News 5 922 231 288
SciERC  (Luan et al., 2018) Scientific 7 1366 187 397
Semeval-RE  (Hendrickx et al., 2010) Scientific 10 6478 1492 2714
NYT11-HRL  (Takanobu et al., 2019) News 12 60765 146 362
KBP37  (Zhang and Wang, 2015) News 18 15911 1723 3405
NYT  (Riedel et al., 2010) News 24 54412 4975 4985
Wiki-ZSL  (Chen and Li, 2021) $\dagger$ Wikipedia 83 - - -
FewRel  (Han et al., 2018) $\dagger$ Wikipedia 100 - - -
RE-zh CMeIE  (Guan et al., 2020) Biomedical 53 14339 3585 -
DuIE2.0  (Li et al., 2019) News 49 171126 20652 -
COAE2016$\dagger$ 5 General 9 - - 971
[…]
SKE2020$\dagger$ 6 News 49 - - 3601
Table 10: Statistical data of Relation Extraction (RE) datasets, with an asterisk ($\dagger$) indicating a zero-shot evaluation set not included in the instruction tuning. The test sets for CMeIE and DuIE2.0 are not open-sourced, thus we use the validation sets as our evaluation data. For the FewRel and Wiki-ZSL datasets, we follow  Chia et al. (2022).
Task Dataset Domain #Schemas #Train #Val #Test
EE-en ACE2005  (Walker et al., 2006) News 33(22) 3257 319 293
CASIE  (Satyapanich et al., 2020) Cybersecurity 5(26) 3732 777 1492
PHEE  (Sun et al., 2022) Biomedical 2(16) 2897 960 968
CrudeOilNews $\dagger$  (Lee et al., 2022b) Oil News 18(104) - - 356
RAMS $\dagger$  (Ebner et al., 2020) News 106(398) - - 887
WikiEvents $\dagger$  (Li et al., 2021) Wikipedia 31(81) - - 249
EE-zh DuEE1.0  (Li et al., 2020b) News 65(217) 11908 1492 -
DuEE-Fin  (Han et al., 2022) Finance 13(91) 7015 1171 -
FewFC $\dagger$  (Zhou et al., 2021) Finance 5(29) - - 2879
CCF law $\dagger$7 Law 9(39) - - 971
Table 11: Statistical data of Event Extraction (EE) datasets, with an asterisk ($\dagger$) indicating a zero-shot evaluation set not included in the instruction tuning. The test sets for DuEE1.0 and DuEE-Fin are not open-sourced, thus we use the validation sets as our evaluation data.

### v3

@@
 We construct IEPile by collecting and cleaning 33 existing IE datasets, and introduce schema-based instruction generation to unearth a large-scale corpus.
-Experimental results on LLaMA, Baichuan and Qwen demonstrate that using IEPile can enhance the performance of LLMs for IE, especially the zero-shot generalization.
+Experimentally, IEPile enhance the performance of LLMs for IE, with notable improvements in zero-shot generalization.
 We open-source the resource and pre-trained models, hoping to provide valuable support to the NLP community.
@@
 Information Extraction Corpus
-Honghao ${}^{\spadesuit}$${}^{\diamondsuit}$thanks:     Equal Contribution., Lin ${}^{\clubsuit}$${}^{\diamondsuit}$11footnotemark: 1, Hongbin ${}^{\spadesuit}$, Ningyu ${}^{\spadesuit}$${}^{\diamondsuit}$thanks:     Corresponding Author.
-Mengshu ${}^{\clubsuit}$${}^{\diamondsuit}$, Lei ${}^{\clubsuit}$${}^{\diamondsuit}$, Huajun ${}^{\spadesuit}$${}^{\diamondsuit}$22footnotemark: 2 ${}^{\spadesuit}$ Zhejiang University ${}^{\diamondsuit}$ Zhejiang University - Ant Group Joint Laboratory of Knowledge Graph ${}^{\clubsuit}$ Ant Group {guihonghao,zhangningyu,huajunsir}@zju.edu.cn https://github.com/zjunlp/IEPile
+Honghao Guithanks:     Equal Contribution., Lin Yuan11footnotemark: 1, Hongbin Ye, Ningyu Zhangthanks:     Corresponding Author.
+Mengshu Sun, Lei Liang, Huajun Chen22footnotemark: 2 Zhejiang University Zhejiang University - Ant Group Joint Laboratory of Knowledge Graph Ant Group {guihonghao,zhangningyu,huajunsir}@zju.edu.cn https://github.com/zjunlp/IEPile
 1 Introduction
@@
 Finally, we obtain IEPile which contains approximately 0.32B tokens.
-By fine-tuning the Baichuan (Yang et al., 2023), LLaMA (Touvron et al., 2023b), and Qwen (Bai et al., 2023), we show that LLMs with IEPile can yield better zero-shot performance than baselines.
+By fine-tuning a selection of the latest prominent models (Yang et al., 2023; Touvron et al., 2023b; Bai et al., 2023) on the IEPile dataset, we show that LLMs with IEPile can yield better zero-shot performance than baselines.
 This achievement not only verifies the effectiveness of the IEPile dataset but also provides a framework for creating IE datasets in other domains.
@@
 To broadly cover various domains and meet the practical demands, we collect datasets necessary for IE from multiple data sources.
-Our corpus mainly involves bilingual data (Chinese and English) and focuses on three main categories of IE tasks: Named Entity Recognition (NER), Relation Extraction (RE), and Event Extraction (EE).
+Our corpus mainly involves bilingual data (Chinese and English) and focuses on three principal categories of IE tasks: Named Entity Recognition (NER), Relation Extraction (RE), and Event Extraction (EE).
 In total, we gather 26 English datasets and 7 Chinese datasets.
@@
 2.2 Schema-Based Instruction Generation
-We concentrate on instruction-based information extraction (IE), which relies on three elements to compose an instruction: 1) Task Description, a template utilized to distinguish between different IE tasks; 2) Input Text, the source text that the model is expected to extract information; and 3) Schema sequence, which defines the information that the model is supposed to extract, including entity types, relations, events, etc.
+We concentrate on instruction-based information extraction (IE), a methodology that incorporates three crucial elements to compose an instruction: 1) Task Description, a template utilized to distinguish between different IE tasks; 2) Input Text, the source text to be extracted; and 3) Schema sequence, which defines the information that the model is supposed to extract, including entity types, relations, events, etc.
 Among these, the schema sequence is critical as it reflects the specific extraction requirements and is dynamically variable.
-Therefore, constructing the schema sequence in an instruction is of utmost importance.
+Therefore, the construction of the schema sequence within an instruction holds critical significance.
 Positive and Negative Schema Mechanism in Instructions.
@@
 However, in the era of generative IE, represented by models like UIE  (Lu et al., 2022a), introduce the concept of integrating a schema sequence (refers to as Structural Schema Instructor, or SSI) in the model’s input to guide its output, restricting the range of output to the SSI.
-The concept necessitates including the entire predefined label set of a dataset as the SSI to guide the model’s output during inference.
-Consequently, if the SSI during the training contains only positive schemas, the model will tend to generate corresponding answers for every label within the SSI during inference.
+The method necessitates including the entire predefined label set of a dataset as the SSI to guide the model’s output during inference.
+As a result, if the SSI during the training contains only positive schemas, the model will tend to generate corresponding answers for every label within the SSI during inference.
 Therefore, to make the model explicitly reject generating outputs for negative schemas, it is necessary to incorporate negative schemas into the SSI.
@@
 In traditional approaches, the $Neg\_L$ is simply defined as $L-Pos\_L$.
-However, they overlook a critical aspect: it is important to pay special attention to negative schemas that are semantically close to positive schemas.
+However, they overlook a critical aspect: it is important to pay special attention to negative schemas that are semantically similar to positive schemas.
 Inspired by the theory of contrastive learning, we construct a hard negative schema dictionary $\mathcal{K}$, where each key represents a unique schema and the associated value is a collection of schemas that are semantically similar to the key schema.
@@
 Consequently, even if the number of schemas inquired during the evaluation phase differs from that of training, the batched mechanism allows us to distribute the inquiries across $split\_num$ schemas, thereby mitigating the decline in generalization performance.
-Figure 2: Distribution of different tasks, domains, and datasets within IEPile.
+Figure 2: Distribution of different tasks, domains, and source datasets within the IEPile.
 2.3 Data Statistics
-Based on the aforementioned methods, we obtain IEPile, which includes roughly 2 million instruction entries and approximately 0.32B tokens (utilizing the Baichuan2 tokenizer).
-Figure 2 displays the distribution of domains and datasets within IEPile, including 33 datasets spanning multiple domains such as general, news, finance, and biomedical.
-Additionally, Table B.2 provides examples of instructions and outputs for 3 different tasks within the IEPile dataset.
+Based on the aforementioned methods, we obtain the IEPile dataset, which includes roughly 2 million instruction entries and approximately 0.32B tokens (utilizing the Baichuan2 tokenizer).
+Figure 2 displays the distribution of domains and source datasets within the IEPile, including 33 datasets spanning multiple domains such as general, news, finance, and biomedical.
+Additionally, Table C.6 provides examples of instructions and outputs for 3 different tasks within the IEPile.
 Method NER RE EE
@@
 Qwen1.5-IEPile 57.90 40.92 38.49 39.71 11.38 21.26 30.69 21.11
+LLaMA3-IEPile 56.11 35.58 37.18 36.38 9.71 20.27 39.88 23.29
+OneKE 60.91 39.19 42.18 40.68 12.43 22.58 38.49 24.50
 Table 1: Zero-shot performance on English datasets.
@@
 For the task of event extraction, we only present the results of event detection in the main text.
+3 Experiments
+Based on IEPile, we fine-tune several latest prominent models, then compare their zero-shot generalization capabilities against a range of baseline models.
+Results of the full supervision evaluation and training details are described in Appendix C.
+3.1 Experimental Settings
+Evaluation Metrics: We employ span-based Micro-F1 as the metric for measuring model performance.
+Baselines: We select a range of strong models for comparative analysis, which include UIE, Baichuan2-13B-Chat, GPT-4, InstructUIE, YAYI-UIE, and so on.
+Zero-shot Benchmark: We collect 13 datasets that are not present in the training set.
+OneKE: Additionally, we perform full-parameter fine-tuning of the alpaca2-chinese-13b model utilizing IEPile and other proprietary information extraction datasets.
+3.2 Main Results
+In Tables 1 and  2, we report the zero-shot performance across three tasks and two languages.
+Overall, after training with the IEPile, the models achieve better results in the majority of tasks.
+We believe the success is due to the hard negative schema construction and batched instruction generation strategy, which can mitigate the train-eval mismatch and semantic ambiguity for the diverse schema.
+We also observe that IEPile-models are slightly behind GPT-4 in English NER.
+We hypothesize that the marginal gap may be attributed to GPT-4’s exposure to a vast corpus of similar data during its training.
+Moreover, it is essential to note that InstructUIE focuses on English data while IEPile incorporates both English and Chinese data.
+This disparity in data may influence the capability of the model in English, potentially reducing the performance.
+Additionally, OneKE achieves the best results in nearly all zero-shot evaluation tasks.
+We attribute this success to the enhancements brought by full parameter fine-tuning.
+3.3 Analysis
+Inconsistency in the Number of Schema Queries Hurt Generalization.
+We investigate the impact on model performance when different numbers of schema queries are used during the training and evaluation.
+We train the Baichuan2 using full-schema instructions on 3 datasets: Ontonotes (18 schemas), DuIE2.0 (49 schemas), and ACE2005 (33 schemas).
+For the evaluation, we test the model using two strategies: one with the full set of schema queries and another with a fixed set of 10 schema queries.
+The results depicted in Figure 3 (a) indicate that the mismatch in the number of schema queries during evaluation significantly reduces the model’s performance.
+Further analysis of the model’s outputs reveals that the model always tends to generate outputs for each inquiry.
+We hypothesize that the number of schema queries is one of the key factors affecting the generalization ability.
+The model needs to first adapt to the number of schema inquiries that are rare during the training and then adapt to the unseen schema.
 Method NER RE EE
@@
 Qwen1.5-IEPile 63.08 37.50 50.29 72.29 50.70 30.55 51.18 78.77 61.43 70.10
+LLaMA3-IEPile 61.88 37.43 49.66 73.67 48.12 31.29 51.03 81.52 59.92 70.72
+OneKE 72.61 35.06 53.84 74.15 49.83 29.95 51.31 80.11 62.19 71.15
 Table 2: Zero-shot performance on Chinese datasets. Since UIE and InstructUIE do not train with Chinese data, we do not report performance of these two models on Chinese datasets.
-3 Experiments
-Based on IEPile, we fine-tune (Lora) models such as Baichuan2, LLaMA2, and Qwen1.5, subsequently comparing their zero-shot generalization capabilities against a range of baseline models.
-Results of the full supervision evaluation, and training details are in Appendix C.
-3.1 Experimental Settings
-Evaluation Metrics: We employ span-based Micro-F1 as the metric for measuring model performance.
-Baselines: We select a range of strong models for comparative analysis, which include UIE  (Lu et al., 2022a), LLaMA2-13B-Chat, Baichuan2-13B-Chat, Mistral-7B-Instruct-v0.2  (Jiang et al., 2023), Qwen1.5-14B-Chat, ChatGPT  (Ouyang et al., 2022), GPT-4  (OpenAI, 2023), InstructUIE  (Wang et al., 2023b), and YAYI-UIE  (Xiao et al., 2023).
-Zero-shot Benchmark: We collect 13 datasets that are not present in the training set.
-3.2 Main Results
-In Tables 1 and  2, we report the zero-shot performance across three tasks: NER, RE, and EE for two languages.
-Overall, after training with the IEPile, the models achieve better results in the majority of tasks.
-We believe the success is due to the hard negative schema construction and batched instruction generation strategy, which can mitigate the train-eval mismatch and semantic ambiguity for the diverse schema.
-We also observe that Baichuan2-IEPile, LLaMA2-IEPile and Qwen1.5-IEPile are slightly behind GPT-4 in English NER.
-We hypothesize that the marginal gap may be attributed to GPT-4’s exposure to a vast corpus of similar data during its training.
-Moreover, it is essential to note that InstructUIE focuses on English data while IEPile incorporates both English and Chinese data.
-This disparity in data may influence the capability of the model in English, potentially reducing the performance.
-Inconsistency in the Number of Schema Queries Hurt Generalization.
-We investigate the impact on model performance when different numbers of schema queries are used during the training and evaluation.
-We train the Baichuan2 using full-schema instructions on 3 datasets: Ontonotes (18 schemas), DuIE2.0 (49 schemas), and ACE2005 (33 schemas).
-For the evaluation, we test the model using two strategies: one with the full set of schema queries and another with a fixed set of 10 schema queries.
-The results depicted in Figure 3(a) indicate that the mismatch in the number of schema queries during evaluation significantly reduces the model’s performance.
-Further analysis of the model’s outputs reveals that the model always tends to generate outputs for each inquiry.
-We hypothesize that the number of schema queries is one of the key factors affecting the generalization ability.
-The model needs to first adapt to the number of schema inquiries that are rare during the training and then adapt to the unseen schema.
-Figure 3: (a) When there is an inconsistency in the number of schema inquiries during the training and evaluation, the performance of the model significantly decreases.
-(b) The impact of removing the hard negative schema dictionary on the performance of the model.
-3.3 Analysis
 Inadequate Differentiation Among Schemas Lead to Semantic Similar Confusion.
@@
 Therefore, processing instructions that are semantically prone to confusion poses significant challenges, and the hard negative schema dictionary plays a crucial role in bolstering model robustness and improving the accuracy of predictions.
+Figure 3: (a) When there is an inconsistency in the number of schema inquiries during the training and evaluation, the performance of the model significantly decreases.
+(b) The impact of removing the hard negative schema dictionary on the performance of the model.
 4 Conclusion and Future Work
@@
 Besides, IEPile is confined to data in English and Chinese, and in the future, we hope to include data in more languages.
-From the model’s perspective, our research evaluates only two models, Baichuan and LLaMA, along with a few baselines due to the computation resources.
-IEPile can be applied to any other LLMs such as ChatGLM Du et al.
+From the model’s perspective, our research evaluates limited models, along with a few baselines due to the computation resources.
+Theoretically, IEPile can be applied to any other LLMs such as ChatGLM Du et al.
 (2022) and Gemma  [^3] https://huggingface.co/google/gemma-7b..
@@
 We conduct a thorough review of the data used, verifying the legitimacy of their sources and compliance with their usage, thus avoiding any infringement on personal privacy or involvement with unauthorized information.
+Acknowledgements
+We would like to express our sincere gratitude to the anonymous reviewers for their thoughtful and constructive feedback.
+This work was supported by the National Natural Science Foundation of China (No.
+62206246), the Fundamental Research Funds for the Central Universities (226-2023-00138), Zhejiang Provincial Natural Science Foundation of China (No.
+LGG22F030011), Yongjiang Talent Introduction Programme (2021A-156-G), CCF-Baidu Open Fund, and Information Technology Center and State Key Lab of CAD&CG, Zhejiang University.
 Appendix A Related Work
 A.1 Information Extraction Datasets
-Large-scale pre-trained corpora are crucial for the effectiveness of LLMs, providing a wealth of knowledge and a foundation for language comprehension, while the annotated data for information extraction (IE) also holds its importance.
-Although the field of IE has accumulated a considerable amount of annotated data (Walker et al., 2006; Riedel et al., 2010; Sang and Meulder, 2003; Luan et al., 2018), these datasets are often limited in size, scattered in distribution, and lack standardization in schema.
-Faced with these limitations, there is an urgent need for generating instruction data through unified and automated methods to fill the gap of the current lack of centralized, large-scale IE instruction tuning datasets.
-We focus on instruction-based IE scenarios and, by collecting and cleaning existing IE data, build a large, schema-rich IE instruction dataset called IEPile.
+Large-scale pre-trained corpora are crucial for the effectiveness of LLMs, providing a wealth of knowledge and a foundation for language comprehension.
+At the same time, the annotated data for information extraction (IE) also holds its importance.
+Although the field of IE has accumulated a considerable amount of annotated data (Walker et al., 2006; Riedel et al., 2010; Sang and Meulder, 2003; Luan et al., 2018; Gui et al., 2023), these datasets are often limited in size, scattered in distribution, and lack standardization in schema.
+Faced with these limitations, there is an urgent need for generating instruction data through unified and automated methods to bridge the gap presented by the current absence of centralized, large-scale IE instruction datasets.
+In this paper, we concentrate on instruction-based IE scenarios.
+We develop a comprehensive, schema-rich instruction dataset for IE by collecting and cleaning existing IE datasets, called IEPile.
 IEPile is designed to enhance the adaptability and processing capabilities of LLMs for different IE tasks, simultaneously strengthening their generalization skills to extract from new domains and schemas.
@@
 Recently, LLMs (Brown et al., 2020; Ouyang et al., 2022; Touvron et al., 2023a, b) demonstrate their exceptional versatility and generalization capabilities across a variety of downstream tasks  (Vilar et al., 2023; Hegselmann et al., 2023).
-Particularly in the domain of IE, these models have the potential to tackle many challenges previously encountered in research  (Zheng et al., 2017; Li et al., 2020a; Paolini et al., 2021; Lu et al., 2022b; Lou et al., 2023; Chen et al., 2022b, 2024), such as adaptability issues when dealing with unseen labels during training.
+Particularly in the domain of IE, these models have the potential to tackle many challenges previously encountered in research  (Zheng et al., 2017; Li et al., 2020a; Paolini et al., 2021; Lu et al., 2022b; Lou et al., 2023; Chen et al., 2022b, 2024), such as adaptability issues when dealing with unseen labels.
 Some studies  (Wei et al., 2023; Wang et al., 2023a; Xie et al., 2023) make significant performance gains in low-resource settings by designing prompt-based frameworks and leveraging models like ChatGPT for in-context learning.
@@
 Additional research explore areas such as prompt learning  (Zhang et al., 2023a), guidelines  (Sainz et al., 2023) and synthetic dataset  (Amalvy et al., 2023).
-Despite these advancements, current models fine-tuned with an instruction approach face a major challenge: the coarse schema handling strategies in constructing instructions can lead to a mismatch in the number of schema queries between the training and evaluation phases, thus weakening the model’s generalizability.
+Despite these advancements, current models fine-tuned with instruction data face a major challenge: the coarse schema handling strategies in constructing instructions could potentially impair the models’ capacity for generalization.
 Appendix B Construction Details of IEPile
@@
 Data Collection
-To comprehensively cover various domains and meet the practical demands of IE, we collect IE datasets from multiple sources.
-IEPile dataset mainly involves bilingual data (Chinese and English) and focuses on three main categories of IE tasks: Named Entity Recognition (NER), Relation Extraction (RE), and Event Extraction (EE).
-The English data mainly comes from the benchmark dataset IEINSTRUCTIONS  (Wang et al., 2023b), while the Chinese data is similar to the Chinese datasets mentioned in the YAYI-UIE  (Xiao et al., 2023).
+To comprehensively cover various domains and meet the practical demands of information extraction (IE), we collect IE datasets from multiple sources.
+IEPile dataset mainly involves bilingual data (Chinese and English) and three IE tasks: Named Entity Recognition (NER), Relation Extraction (RE), and Event Extraction (EE).
+The English part mainly comes from the benchmark dataset IEINSTRUCTIONS  (Wang et al., 2023b), while the Chinese data is similar to the Chinese datasets mentioned in the YAYI-UIE  (Xiao et al., 2023).
 It should be noted that our Chinese dataset collection is conducted concurrently with the aforementioned research.
@@
 If texts from the test set appear previously in the training or validation sets, we would exclude these instances from the training and validation sets.
-Moreover, we formulate three heuristic rules to eliminate low-quality and meaningless data:
+Furthermore, we formulate three heuristic rules to eliminate low-quality and meaningless data:
 1) Non-alphabetic characters comprising more than 80% of the text;
@@
 The rationale behind the development of these hard negatives is two-fold: firstly, to induce a more frequent co-occur of semantically similar schemas within the instructions, and secondly, to reduce the volume of training instances without sacrificing the model’s performance.
-Consider a scenario involving 48 schemas paired with a $split\_num$ of 4.
-Conventionally, this setup would necessitate the generation of 12 distinct instructions.
-Yet, with the inclusion of hard negatives, this number can be reduced to as few as 3.
+In the context of a dataset comprising 48 schemas with a given $split\_num$ of 4, traditional mode would dictate the creation of 12 unique instructions per data point.
+However, through the integration of hard negatives, this requisite can be substantially minimized to a mere 3 instructions.
 Algorithm 1 Schema-Based Instruction Generation
-0:  Text $S$, Full label set $L$, Positive label set $Pos\_L$, Number of schemas to split $split\_num$
-0:  Negative label set $Neg\_L$, Set of instructions $Instructions$ Step 1: Initialize Hard Negative Schema Dictionary $\mathcal{K}$
+0:  Text $S$, Predefined label set $L$, Positive schema set $Pos\_L$, Number of schemas to split $split\_num$
+0:  Set of $Instructions$ Step 1: Initialize Hard Negative Schema Dictionary $\mathcal{K}$
   for all schema in $L$ do
-     $\mathcal{K}[schema]\leftarrow\text{SEMANTIC-CLOSE}(schema,L)$
+     $\mathcal{K}[schema]\leftarrow\text{SEMANTIC-SIMILAR}(schema,L)$
   end for
@@
   $Neg\_L\leftarrow Hard\_L\cup Other\_L$
-  $L^{\prime}=Neg\_L\cup Pos\_L$
-  Shuffle $L^{\prime}$ to get a randomized sequence
+  $L^{\prime}\leftarrow Neg\_L\cup Pos\_L$
+  Shuffle $L^{\prime}$ to obtain a randomized sequence
   Step 3: Batched Instruction Generation
@@
 Selection of split_num.
-In determining the range for $split\_num$, we combine empirical experimental results with an in-depth analysis of dataset characteristics.
+In the determination of the optimal range for $split\_num$, our methodology integrates empirical results with an in-depth analysis of dataset characteristics.
 For a dataset containing N different labels, the theoretical value of $split\_num$ should fall within the interval [1, N].
-Given that we deal with datasets of varying numbers of labels, we aim to identify a $split\_num$ value that could be shared by as many datasets as possible, thereby ensuring that this number is a common divisor for the label numbers of most datasets.
-For instance, for Named Entity Recognition datasets, we set $split\_num$ to 6; for Relation Extraction and Event Extraction datasets, we establish $split_{n}um$ at 4.
+Addressing datasets with heterogeneous label counts, our objective is to identify a $split\_num$ value that offers broad applicability across numerous datasets, thus ensuring this value serves as a common divisor for the majority of dataset label counts.
+For instance, for Named Entity Recognition datasets, we set $split\_num$ to 6; for Relation Extraction and Event Extraction datasets, we establish $split\_num$ at 4.
 We also observe that when $split\_num$ is 1, the ratio of positive to negative samples significantly impacts model performance, and the corresponding number of training samples becomes vast, affecting efficiency adversely.
 More crucially, we believe that enumerating multiple schemas in instructions aids the model in more effectively learning to distinguish and identify various schemas, thereby enhancing model performance.
-Moreover, to enhance model robustness and its clear understanding of the dynamically changing schema sequences in instructions, we set the actual number of schema splits within a dynamic range of [$split\_num$ // 2, $split\_num$ + $split\_num$ // 2].
-Specifically, if the number of schemas in the last batch is less than half of $split\_num$, it is merged with the previous batch; it is only treated as a separate batch when its number exceeds half of $split\_num$.
-Task Instruction & Output
-NER
-1{
-2    "instruction": "You are an expert in named entity recognition.
-Please extract entities that match the schema definition from the input.
-Return an empty list if the entity type does not exist.
-Please respond in the format of a JSON string.",
-3    "schema": ["location", "else", "organization", "person"],
-4    "input": "The objective of the Basic Course on War is to provide for combatants of the EPR basic military knowledge for the armed conflict against the police and military apparatus of the bourgeoisie."
-5}
-6output = {
-7    "location": [],
-8    "else": [],
-9    "organization": ["EPR"],
-10    "person": []
-11}
-RE
-1{
-2    "instruction": "You are an expert in relationship extraction.
-Please extract relationship triples that match the schema definition from the input.
-Return an empty list for relationships that do not exist.
-Please respond in the format of a JSON string.",
-3    "schema": ["place of birth", "country capital", "country of administrative divisions", "company"],
-4    "input": "Born on May 1 , 1927 , in Brichevo , Bessarabia in the present-day Republic of Moldova , Mr.
-Bertini emigrated to Palestine with his family as a child and pursued musical studies there , in Milan , and in Paris , where he worked with Nadia Boulanger and Arthur Honegger."
-5}
-6output = {
-7    "place of birth": [{"head": "Mr. Bertini", "tail": "Paris"}],
-8    "country capital": [],
-9    "country of administrative divisions": [],
-10    "company": []
-11}
-EE
-1{
-2    "instruction": "You are an expert in event extraction.
-Please extract events from the input that conform to the schema definition.
-Return an empty list for events that do not exist, and return NAN for arguments that do not exist.
-If an argument has multiple values, please return a list.
-Respond in the format of a JSON string.",
-3    "schema": [{"event_type": "pardon", "trigger": true, "arguments": ["defendant"]}, {"event_type": "extradite", "trigger": true, "arguments": ["person", "agent", "destination", "origin"]}, {"event_type": "sue", "trigger": true, "arguments": ["place", "plaintiff"]}, {"event_type": "start position", "trigger": true, "arguments": ["person", "entity", "place"]}],
-4    "input": "Ethical and legal issues in hiring Marinello"
-5}
-6output = {
-7    "pardon": [],
-8    "extradite": [],
-9    "sue": [],
-10    "start position": [{"trigger": "hiring", "arguments": {"person": "Marinello", "entity": "NAN", "place": "NAN"}}]
-11}
-Table 3: Instructions and outputs for 3 tasks: Named Entity Recognition (NER), Relation Extraction (RE), and Event Extraction (EE).
-The instruction and output formats for IEPile adopt a structure similar to JSON strings.
+Furthermore, to enhance model robustness and its clear understanding of the dynamically changing schema sequences in instructions, we set the actual number of schema splits within a dynamic range of [$split\_num$ // 2, $split\_num$ + $split\_num$ // 2].
+Specifically, if the number of schemas in the last batch is less than half of $split\_num$, it is merged with the previous batch; otherwise, it stands as an independent batch.
 Instruction Format
 The instruction format of IEPile adopts a structure akin to JSON strings, essentially constituting a dictionary-type string.
-This structure is comprised of three main components: (1) “instruction”, which is the task description outlining the objective of the instruction’s execution; (2) “schema”, a list of labels that need to be extracted, specifying the key fields of the information to be extracted; (3) “input”, the source text from which information is to be extracted.
-Examples of instructions corresponding to various tasks can be found in Table B.2.
+This structure is comprised of three main components: (1) “instruction”, which is the task description outlining the objective of the instruction’s execution; (2) “schema”, a list of labels that need to be extracted; (3) “input”, the source text from which information is to be extracted.
+Examples of instructions corresponding to various tasks can be found in Table C.6.
 B.3 Data Statistics
 Based on the aforementioned methodologies, we construct a high-quality information extraction instruction dataset known as IEPile.
-This dataset contains approximately two million instances, each comprising two fields: “instruction” and “output”, which are formatted for direct use in the supervised fine-tuning of models.
-In terms of storage, IEPile occupies around 3GB of disk space, encompassing roughly 0.32B tokens (utilizing the BaiChuan2 tokenizer).
+This dataset contains approximately two million instances and approximately 0.32B tokens.
+Each instance of IEPile comprises two fields: “instruction” and “output”, which are formatted for direct use in the instruction tuning.
 Appendix C Experiments
@@
 Baselines
-For evaluating zero-shot generalization capabilities, we select a range of strong models for comparative analysis:
+To assess the zero-shot generalization capabilities, we select a range of strong models for comparative analysis:
   • UIE  (Lu et al., 2022a): is a unified text-to-structure generation framework that can model various information extraction (IE) tasks generically.
@@
   • ChatGPT  (Ouyang et al., 2022): also known as GPT-3.5-turbo, represents the most advanced artificial intelligence language model with chat optimization capabilities to date.
+  • GPT-4  (OpenAI, 2023): Known as the most powerful closed-source chat model to date.
+  • LLaMA3  [^4] https://ai.meta.com/blog/meta-llama-3/.: The latest release in the LLaMA model series, achieving significant improvements across various benchmarks.
   • InstructUIE  (Wang et al., 2023b): a unified IE framework based on multi-task instruction tuning.
   • YAYI-UIE  (Xiao et al., 2023): is an end-to-end, chat-enhanced, universal information extraction framework that supports both Chinese and English, fine-tuned with instructional prompts for generalized information.
-Zero-shot Dataset
-To ensure the validity of the zero-shot evaluation and prevent result bias due to data similarity, we select datasets primarily derived from news and biomedical fields as our training sets.
-This selection is intended to train the model’s capability for instruction following and schema-based extraction.
-For the evaluation data, we adopt the 13 cross-domain datasets recommended in IEINSTRUCTIONS and YAYI-UIE, which include: for Named Entity Recognition (NER) tasks, we use the CrossNER  (Liu et al., 2021), Weibo NER  (Peng and Dredze, 2015), and Boson [^4] https://github.com/InsaneLife/; in Relation Extraction (RE) tasks, we involve FewRel  (Han et al., 2018), Wiki-ZSL  (Chen and Li, 2021), COAE2016555https://github.com/Sewens/COAE2016, IPRE  (Wang et al., 2019), and SKE2020666https://aistudio.baidu.com/datasetdetail/177191; and for Event Extraction (EE), we include RAMS  (Ebner et al., 2020), WikiEvents  (Li et al., 2021), CrudeOilNews  (Lee et al., 2022b), FewFC  (Zhou et al., 2021), and CCF law  [^7] https://aistudio.baidu.com/projectdetail/4201483.
-These datasets cover a wide range of fields including literature, music, law, and oil news.
-It is noteworthy that these evaluation data sets are not used during the training, ensuring that our evaluation accurately reflects the model’s generalization and adaptation capabilities for unseen domains and unseen schema data in zero-shot information extraction.
-Hyperparameter Value
-Number of Epochs 5
-Learning Rate 5e-5
-Batch Size 20
-Accumulate 4
-Optimizer Adam
-Lora_r 64
-Lora_alpha 64
-Lora_dropout 0.05
-Table 4: Training Hyperparameters
+C.2 OneKE
+We leverage IEPile, InstructIE (Gui et al., 2023), CMRC (Cui et al., 2019), along with certain proprietary business information extraction datasets from Ant Group, to compile a comprehensive training dataset consisting of 2.5 million instances.
+Subsequently, we undertake full-parameter fine-tuning of the alpaca2-chinese-13b [^5] https://huggingface.co/hfl/chinese-alpaca-2-13b.
+model on this training dataset, resulting in the refined model named OneKE.
 Method EN CH
@@
 Qwen1.5-IEPile 11.38 21.26 30.69 21.11 78.77 61.43 70.10
+LLaMA3-IEPile 9.71 20.27 39.88 23.29 81.52 59.92 70.72
+OneKE 12.43 22.58 38.49 24.50 80.11 62.19 71.15
 Argument LLaMA2 0.00 0.00 0.00 0.00 0.00 0.06 0.03
@@
 Qwen1.5-IEPile 11.93 10.57 20.22 14.24 59.49 58.86 59.18
+LLaMA3-IEPile 12.10 10.96 19.20 14.09 48.19 42.59 45.39
+OneKE 11.88 13.26 20.11 15.08 58.83 62.38 60.61
 Table 3: Zero-shot performance on Event Extraction (EE) task. Within each column, shadow and shadow represent the top 2 results.
-C.2 Zero-shot performance on Event Extraction
-As illustrated in Table  5, the model trained with IEPile exhibits outstanding performance in zero-shot event extraction (EE) tasks, surpassing other competing models.
+Zero-shot Dataset
+To ensure the validity of the zero-shot evaluation and prevent result bias due to data similarity, we select datasets primarily derived from news and biomedical fields as our training sets.
+This selection is intended to train the model’s capability for instruction following and schema-based extraction.
+For the evaluation data, we adopt the 13 cross-domain datasets recommended in IEINSTRUCTIONS and YAYI-UIE, which include: for Named Entity Recognition (NER) tasks, we use the CrossNER  (Liu et al., 2021), Weibo NER  (Peng and Dredze, 2015), and Boson [^6] https://github.com/InsaneLife/; in Relation Extraction (RE) tasks, we choose FewRel  (Han et al., 2018), Wiki-ZSL  (Chen and Li, 2021), COAE2016777https://github.com/Sewens/COAE2016, IPRE  (Wang et al., 2019), and SKE2020888https://aistudio.baidu.com/datasetdetail/177191; and for Event Extraction (EE), we include RAMS  (Ebner et al., 2020), WikiEvents  (Li et al., 2021), CrudeOilNews  (Lee et al., 2022b), FewFC  (Zhou et al., 2021), and CCF law  [^9] https://aistudio.baidu.com/projectdetail/4201483.
+These datasets cover a wide range of fields including literature, music, law, and oil news.
+It is noteworthy that these evaluation data sets are not used during the training, ensuring that our evaluation accurately reflects the model’s generalization and adaptation capabilities for unseen domains and unseen schema data in zero-shot information extraction.
+C.3 Zero-shot performance on Event Extraction
+As illustrated in Table  3, the model trained with IEPile exhibits outstanding performance in zero-shot event extraction (EE) tasks, surpassing other baselines.
 Notably, in the Chinese EE task, the LLaMA2-IEPile model’s performance is slightly inferior to YAYI-UIE’s, revealing LLaMA2’s limitations in processing Chinese data.
@@
 This contrast highlights the potential influence of language type on model performance.
-C.3 Hyper-parameter
-In our research, we select two pre-trained models, Baichuan2-13B-Chat and LLaMA2-13B-Chat, as the base models for our study.
+Hyperparameter Value
+Number of Epochs 5
+Learning Rate 5e-5
+Batch Size 20
+Accumulate 4
+Optimizer Adam
+Lora_r 64
+Lora_alpha 64
+Lora_dropout 0.05
+Table 4: Training Hyperparameters
+C.4 Hyper-parameter
+In our research, we select four pre-trained models, Baichuan2-13B-Chat and LLaMA2-13B-Chat, Qwen1.5-14B-Chat, and LLaMA3-8B-Instruct, as the base models for our study.
 Specifically, we employ the LoRA  (Hu et al., 2022) technique and utilize 8 NVIDIA A800 GPUs to perform instruction tuning on our IEPile dataset.
-The training period for each model spans approximately three days.
 Detailed configurations of the hyperparameters during the fine-tuning process are presented in Table 4.
-Dataset InstructUIE YAYI-UIE Baichuan2-IEPile LLaMA2-IEPile
-ACE2005 86.66 81.78 81.86 81.14
-AnatEM 90.89 76.54 87.21 86.90
-BC2GM 85.16 82.05 80.73 83.07
-BC4CHEMD 90.30 88.46 90.45 90.07
-BC5CDR 89.59 83.67 88.07 88.01
-CoNLL2003 92.94 96.77 92.49 92.98
-FabNER 76.20 72.63 77.07 76.33
-FindVehicle 89.47 98.47 98.49 97.91
-GENIA-Ent 74.71 75.21 76.66 77.32
-HarveyNER 88.79 69.57 67.70 62.64
-MIT Movie 89.01 70.14 88.23 89.54
-MIT Restaurant 82.55 79.38 79.85 81.30
-MultiNERD 92.32 88.42 94.60 94.24
-NCBI-Disease 90.23 87.29 85.26 87.59
-Ontonotes 90.19 87.04 87.55 90.34
-Avg 87.27 82.49 85.08 85.29
-MSRA - 95.57 87.99 86.32
-Resume NER - - 93.92 92.86
-CLUE NER - - 80.19 76.57
-Table 6: Overall supervision results on Named Entity Recognition (NER) datasets. Within each row, shadow and shadow represent the top 2 results.
-Dataset InstructUIE YAYI-UIE Baichuan2-IEPile LLaMA2-IEPile
-ADE Corpus 82.31 84.14 83.73 85.87
-CoNLL2004 78.48 79.73 72.87 73.71
-GIDS 81.98 72.36 74.71 74.13
-KBP37 36.14 59.35 65.09 61.49
-NYT 90.47 89.97 93.0 92.22
-NYT11-HRL 56.06 57.53 53.19 54.86
-SciERC 45.15 40.94 43.53 44.58
-Semeval-RE 73.23 61.02 58.47 57.61
-Avg 67.98 68.13 68.07 68.06
-CMeIE - - 49.16 47.40
-DuIE2.0 - 81.19 75.61 74.34
-Table 7: Overall supervision results on Relation Extraction (RE) datasets. Within each row, shadow and shadow represent the top 2 results.
-Dataset InstructUIE YAYI-UIE Baichuan2-IEPile LLaMA-IEPile
-Trigger ACE2005 77.13 65.00 72.46 70.63
-CASIE 67.80 63.00 60.07 61.27
-PHEE 70.14 63.00 66.22 68.52
-Avg 71.69 63.67 66.25 66.81
-DuEE1.0 - 85.00 86.73 84.01
-DuEE-fin - 82.50 83.54 79.00
-Argument ACE2005 72.94 62.71 63.90 62.69
-CASIE 63.53 64.23 56.07 56.78
-PHEE 62.91 77.19 70.85 71.33
-Avg 66.46 68.04 63.60 63.61
-DuEE1.0 - 79.08 75.63 73.79
-DuEE-fin - 70.02 79.34 73.08
-Table 8: Overall supervision results on Event Extraction (EE) datasets. Within each row, shadow and shadow represent the top 2 results.
-C.4 Supervision Results
+Dataset Supervised Zero-shot
+ACE2004 84.28 77.01
+People Daily 98.34 95.29
+Table 5: The results of individual LoRA fine-tuning on ACE2004 and People Daily datasets for Baichuan2-13B-Chat, compared with the zero-shot generalization results of Baichuan2-IEPile on these two datasets.
+C.5 Supervision Results
+Due to limited computational resources, I report only the supervised results for the Baichuan2-IEPile, LLaMA2-IEPile, and OneKE models.
 Tables  6,  7, and  8 present our experimental results under a supervised learning setting on the training dataset.
-Specifically, it can be observed that after training on the IEPile, the model excels in Named Entity Recognition (NER), Relation Extraction (RE), and Event Detection (ED), ranking second across these tasks.
-The model’s performance is only slightly behind other reference models in the Event Argument Extraction.
+Specifically, it can be observed that after training on the IEPile, the model excels in Named Entity Recognition (NER), Relation Extraction (RE), and Event Detection (ED), ranking top 2 across these tasks.
+The model’s performance is only slightly behind other baselines in the Event Argument Extraction.
 Additionally, we record the model’s performance in Chinese NER, RE, and EE tasks, where it demonstrates robust results.
-Notably, unlike fully parameter-fine-tuned models such as InstructUIE and YAYI-UIE, we employ the Low-Rank Adaptation (LoRA) technique for model optimization, which involves partial retraining.
-This difference in fine-tuning approach may have a certain impact on the results of supervision learning.
 In a comprehensive assessment, the IEPile-trained model showcases performance on par with other models in instruction-based information extraction (IE) tasks and significantly improves performance in zero-shot IE tasks compared to other models.
 This indicates the significant application prospects and potential of IEPile in the current field of IE.
-C.5 Impact of Potential Dataset Bias on Model Performance and Generalization
-During the IEPile project, we identify that potential biases introduced by the datasets used can affect the model’s performance and generalization capability.
+C.6 Impact of Potential Dataset Bias on Model Performance and Generalization
+During the research, we identify that potential biases introduced by the datasets used can affect the model’s performance and generalization capability.
 Firstly, biases in the definition of schemas within the datasets have a negative impact on model performance  (Huang et al., 2024).
-In the early stages of training, we observe instability in results due to mutual interference among multiple datasets that contain the same schemas but with differing definitions (such as wikiann, wikineural, polyglot-NER, and CoNLL2003).
-Consequently, in the later stages, we remove all such datasets except for CoNLL2003.
+In the early stages of training, we observe instability in results due to mutual interference among multiple datasets that contain the same schemas but with differing definitions.
+For instance, despite wikiann, wikineural, polyglot-NER, and CoNLL2003 all containing common schemas such as people and organization, they each possess distinct scheme definitions.
+Consequently, in the later stages, only CoNLL2003 is retained.
 Secondly, the model demonstrates good generalization when dealing with datasets having schemas similar to those in the training set.
-For instance, the model proficiently handles People Daily and ACE2004 datasets, which share similar schemas to the MASR and ACE2005 NER datasets in the training set, despite not being included in the training corpus.
-Lastly, the use of common, coarse-grained labels (such as “person” and “organization”) within the IEPile project lead the model, after training, to favor these coarse categories over fine-grained ones (such as “scientist” and “company”) when predicting instructions that included both levels of granularity.
+As shown in Table 5, despite not being included in the training corpus, the People Daily and ACE2004 NER datasets share similar schemas with the MASR and ACE2005 NER dataset in the training set, and the Baichuan2-IEPile model is still capable of handling them proficiently.
+Lastly, the use of common, coarse-grained labels (such as “person” and “organization”) within the IEPile lead the model, after training, to favor these coarse categories over fine-grained ones (such as “scientist” and “company”) when predicting instructions that included both levels of granularity.
+Dataset InstructUIE YAYI-UIE Baichuan2-IEPile LLaMA2-IEPile OneKE
+ACE2005 86.66 81.78 81.86 81.14 83.45
+AnatEM 90.89 76.54 87.21 86.90 87.88
+BC2GM 85.16 82.05 80.73 83.07 82.05
+BC4CHEMD 90.30 88.46 90.45 90.07 90.56
+BC5CDR 89.59 83.67 88.07 88.01 88.45
+CoNLL2003 92.94 96.77 92.49 92.98 93.04
+FabNER 76.20 72.63 77.07 76.33 81.06
+FindVehicle 89.47 98.47 98.49 97.91 99.45
+GENIA-Ent 74.71 75.21 76.66 77.32 78.29
+HarveyNER 88.79 69.57 67.70 62.64 69.87
+MIT Movie 89.01 70.14 88.23 89.54 89.96
+MIT Restaurant 82.55 79.38 79.85 81.30 79.89
+MultiNERD 92.32 88.42 94.60 94.24 94.69
+NCBI-Disease 90.23 87.29 85.26 87.59 86.95
+Ontonotes 90.19 87.04 87.55 90.34 89.08
+Avg 87.27 82.49 85.08 85.29 86.24
+MSRA - 95.57 87.99 86.32 89.02
+Resume NER - - 93.92 92.86 95.84
+CLUE NER - - 80.19 76.57 78.43
+Table 6: Overall supervision results on Named Entity Recognition (NER) datasets. Within each row, shadow and shadow represent the top 2 results.
+Dataset InstructUIE YAYI-UIE Baichuan2-IEPile LLaMA2-IEPile OneKE
+ADE Corpus 82.31 84.14 83.73 85.87 87.24
+CoNLL2004 78.48 79.73 72.87 73.71 76.16
+GIDS 81.98 72.36 74.71 74.13 76.69
+KBP37 36.14 59.35 65.09 61.49 65.23
+NYT 90.47 89.97 93.00 92.22 94.04
+NYT11-HRL 56.06 57.53 53.19 54.86 55.56
+SciERC 45.15 40.94 43.53 44.58 45.89
+Semeval-RE 73.23 61.02 58.47 57.61 61.46
+Avg 67.98 68.13 68.07 68.06 70.28
+CMeIE - - 49.16 47.40 49.54
+DuIE2.0 - 81.19 75.61 74.34 75.73
+Table 7: Overall supervision results on Relation Extraction (RE) datasets. Within each row, shadow and shadow represent the top 2 results.
+Dataset InstructUIE YAYI-UIE Baichuan2-IEPile LLaMA-IEPile OneKE
+Trigger ACE2005 77.13 65.00 72.46 70.63 71.17
+CASIE 67.80 63.00 60.07 61.27 63.82
+PHEE 70.14 63.00 66.22 68.52 68.60
+Avg 71.69 63.67 66.25 66.81 67.86
+DuEE1.0 - 85.00 86.73 84.01 85.75
+DuEE-fin - 82.50 83.54 79.00 82.91
+Argument ACE2005 72.94 62.71 63.90 62.69 62.75
+CASIE 63.53 64.23 56.07 56.78 57.16
+PHEE 62.91 77.19 70.85 71.33 72.84
+Avg 66.46 68.04 63.60 63.61 64.25
+DuEE1.0 - 79.08 75.63 73.79 75.40
+DuEE-fin - 70.02 79.34 73.08 78.98
+Table 8: Overall supervision results on Event Extraction (EE) datasets. Within each row, shadow and shadow represent the top 2 results.
 Task Dataset Domain #Schemas #Train #Val #Test
@@
 […]
-Table 9: Statistical data of Named Entity Recognition (NER) datasets, with an asterisk ($\dagger$) indicating a zero-shot evaluation set not included in the instruction tuning.
+Table 9: Statistical data of Named Entity Recognition (NER) datasets, with an $\dagger$ indicating the zero-shot evaluation set not included in the training.
 CrossNER  (Liu et al., 2021) is divided into five subsets for our statistical analysis.
-For the zero-shot evaluation sets, we only report their test set.
 Task Dataset Domain #Schemas #Train #Val #Test
@@
 DuIE2.0  (Li et al., 2019) News 49 171126 20652 -
-COAE2016$\dagger$ 5 General 9 - - 971
+COAE2016$\dagger$ 7 General 9 - - 971
 […]
-SKE2020$\dagger$ 6 News 49 - - 3601
-Table 10: Statistical data of Relation Extraction (RE) datasets, with an asterisk ($\dagger$) indicating a zero-shot evaluation set not included in the instruction tuning.
-The test sets for CMeIE and DuIE2.0 are not open-sourced, thus we use the validation sets as our evaluation data.
+SKE2020$\dagger$ 8 News 49 - - 3601
+Table 10: Statistical data of Relation Extraction (RE) datasets, with an $\dagger$ indicating the zero-shot evaluation set not included in the training.
+The test sets for CMeIE and DuIE2.0 are not open-sourced, thus we use the validation sets as our evaluation set.
 For the FewRel and Wiki-ZSL datasets, we follow  Chia et al.
@@
 FewFC $\dagger$  (Zhou et al., 2021) Finance 5(29) - - 2879
-CCF law $\dagger$7 Law 9(39) - - 971
-Table 11: Statistical data of Event Extraction (EE) datasets, with an asterisk ($\dagger$) indicating a zero-shot evaluation set not included in the instruction tuning.
-The test sets for DuEE1.0 and DuEE-Fin are not open-sourced, thus we use the validation sets as our evaluation data.
+CCF law $\dagger$9 Law 9(39) - - 971
+Table 11: Statistical data of Event Extraction (EE) datasets, with an $\dagger$ indicating the zero-shot evaluation set not included in the training.
+The test sets for DuEE1.0 and DuEE-Fin are not open-sourced, thus we use the validation sets as our evaluation set.
+Task Instruction & Output
+NER 1{ 2 "instruction": "You are an expert in named entity recognition.
+Please extract entities that match the schema definition from the input.
+Return an empty list if the entity type does not exist.
+Please respond in the format of a JSON string.", 3 "schema": ["location", "else", "organization", "person"], 4 "input": "The objective of the Basic Course on War is to provide for combatants of the EPR basic military knowledge for the armed conflict against the police and military apparatus of the bourgeoisie." 5} 6output = { 7 "location": [], 8 "else": [], 9 "organization": ["EPR"], 10 "person": [] 11}
+RE 1{ 2 "instruction": "You are an expert in relationship extraction.
+Please extract relationship triples that match the schema definition from the input.
+Return an empty list for relationships that do not exist.
+Please respond in the format of a JSON string.", 3 "schema": ["place of birth", "country capital", "country of administrative divisions", "company"], 4 "input": "Born on May 1 , 1927 , in Brichevo , Bessarabia in the present-day Republic of Moldova , Mr.
+Bertini emigrated to Palestine with his family as a child and pursued musical studies there , in Milan , and in Paris , where he worked with Nadia Boulanger and Arthur Honegger." 5} 6output = { 7 "place of birth": [{"head": "Mr.
+Bertini", "tail": "Paris"}], 8 "country capital": [], 9 "country of administrative divisions": [], 10 "company": [] 11}
+EE 1{ 2 "instruction": "You are an expert in event extraction.
+Please extract events from the input that conform to the schema definition.
+Return an empty list for events that do not exist, and return NAN for arguments that do not exist.
+If an argument has multiple values, please return a list.
+Respond in the format of a JSON string.", 3 "schema": [{"event_type": "pardon", "trigger": true, "arguments": ["defendant"]}, {"event_type": "extradite", "trigger": true, "arguments": ["person", "agent", "destination", "origin"]}, {"event_type": "sue", "trigger": true, "arguments": ["place", "plaintiff"]}, {"event_type": "start position", "trigger": true, "arguments": ["person", "entity", "place"]}], 4 "input": "Ethical and legal issues in hiring Marinello" 5} 6output = { 7 "pardon": [], 8 "extradite": [], 9 "sue": [], 10 "start position": [{"trigger": "hiring", "arguments": {"person": "Marinello", "entity": "NAN", "place": "NAN"}}] 11}
+Table 12: Instructions and outputs for 3 tasks: Named Entity Recognition (NER), Relation Extraction (RE), and Event Extraction (EE).
+The instruction and output formats for IEPile adopt a structure similar to JSON strings.